Roger wrote and solved a system of equations to find the number of nickels (x), dimes (y), and quarters (z) that were in a bank. He found that the general solution was (x,y,z) = ((2x - 4), (8 - z), z). List all of the solutions to Roger's problem.

please show how you reached your answer.
Answer #1

Given . . General Solution was (x,0.2) -((2x-4),(8-2) So, no 21-4 -0. On solving o equation, = = 2-4 2721-* = 4 = In 4 Hence,